OpenStudy (anonymous):

Solve: -2(2x + 5) - 3 = -3(x - 1) 16 -16 -10 -13

OpenStudy (amoodarya):

-2(2x + 5) - 3 = -3(x - 1) -4x-10 -3=-3x +3 -4x+3x =+10 +3 +3 -x=16 so x=-16